# Learning to Summarize by Learning to Quiz: Adversarial Agentic Collaboration for Long Document Summarization

### v2

@@ Abstract
Long document summarization remains a significant challenge for current large language models (LLMs), as existing approaches commonly struggle with information loss, factual inconsistencies, and coherence issues when processing excessively long documents. We propose SummQ, a novel adversarial multi-agent framework that addresses these limitations through collaborative intelligence between specialized agents operating in two complementary domains: summarization and quizzing. Our approach employs summary generators and reviewers that work collaboratively to create and evaluate comprehensive summaries, while quiz generators and reviewers create comprehension questions that serve as continuous quality checks for the summarization process. This adversarial dynamic, enhanced by an examinee agent that validates whether the generated summary contains the information needed to answer the quiz questions, enables iterative refinement through multifaceted feedback mechanisms. We evaluate SummQ on three widely used long document summarization benchmarks. Experimental results demonstrate that our framework significantly outperforms existing state-of-the-art methods across ROUGE and BERTScore metrics, as well as in LLM-as-a-Judge and human evaluations. Our comprehensive analyses reveal the effectiveness of the multi-agent collaboration dynamics, the influence of different agent configurations, and the impact of the quizzing mechanism. This work establishes a new approach for long document summarization that uses adversarial agentic collaboration to improve summarization quality.
@@ 1 Introduction
Summarization has become increasingly critical in modern natural language processing, as organizations and individuals face an ever-growing volume of textual information that requires efficient processing and comprehension (Gambhir & Gupta, 2017; Zhao et al., 2020; Zhang et al., 2021). Prior works in summarization have primarily focused on short to medium-length documents, where models can effectively capture the essential content and generate coherent summaries (See et al., 2017; Fabbri et al., 2019). Recently, there has been an increasing interest in long document document summarization, driven by the need to process extensive texts such as research articles, legal documents, and books (Huang et al., 2021; Kryscinski et al., 2022; Saxena & Keller, 2024).
Recent large language models (LLMs) have shown promising results for summarization tasks (Pu et al., 2023; Laban et al., 2023; Keswani et al., 2024). However, existing methods struggle with long documents, often leading to significant information loss, factual inconsistencies, and difficulty maintaining coherence across lengthy texts (Koh et al., 2023). Current approaches often fail to capture the nuanced relationships between distant parts of a document, resulting in summaries that may miss crucial information or introduce hallucinations (Chrysostomou et al., 2024; Tang et al., 2024a; Xia et al., 2024). Recently, multi-agent systems has demonstrated potential for improving complex reasoning tasks through collaborative interactions (Guo et al., 2024), yet their application to long document summarization remains underexplored (Fang et al., 2024; Kim & Kim, 2025).
@@
To validate the effectiveness of SummQ, we conduct extensive experiments on three long document summarization tasks including MENSA (Saxena & Keller, 2024), BookSum (Kryscinski et al., 2022), and GovReport (Huang et al., 2021). Our results demonstrate that SummQ significantly outperforms existing state-of-the-art methods in terms of ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020), as well as LLM-as-a-Judge and human evaluations. Furthermore, more in-depth analyses highlight the effectiveness of our multi-agent framework in enhancing summary quality, the coverage of the generated quizzes, and the impact of various agent configurations.
@@ 2 Related Work
Recent advances in LLMs have enabled the development of multi-agent systems that harness the strengths of multiple agents to tackle complex tasks (Wang et al., 2024; Guo et al., 2024; Xi et al., 2025). These systems typically involve agent collaboration to boost performance, as seen in multi-agent debating (Du et al., 2024; Xiong et al., 2023; Chen et al., 2023a; Tang et al., 2024b) and discussion (Chen et al., 2024; Saha et al., 2024) for reasoning over short texts (Du et al., 2024; Tang et al., 2024b), paper review (Xu et al., 2023), dataset synthesis (Wang et al., 2025b), machine translation (Wu et al., 2024), and code generation (Huang et al., 2023; Wang et al., 2025a). Collaboration among agents introduces diverse perspectives and complementary skills, leading to higher-quality and more robust outputs.
@@
Long document summarization has seen various methods, including architectural optimizations (e.g., sparse attention (Liu et al., 2021; Ivgi et al., 2023; Bertsch et al., 2023), long-context finetuning (Beltagy et al., 2020; Guo et al., 2021), memory augmentation (Cui & Hu, 2021; Saxena et al., 2025), window extension (Press et al., 2021; Chen et al., 2023b; Su et al., 2024; Yen et al., 2024)) and chunking strategies (e.g., sliding window (Zaheer et al., 2020; Pang et al., 2023). LLMs with improved long-context abilities have shifted the field toward leveraging their strong language skills for summarization (Goyal et al., 2022; Ratner et al., 2023; Keswani et al., 2024), but still face challenges with context limits and maintaining coherence (Pu et al., 2023; Liu et al., 2023). Multi-agent systems have been explored to address these issues, enabling collaborative, more accurate summaries (Zhao et al., 2024; Fang et al., 2024; Jeong et al., 2025), though many still rely on self-verification, leading to biases and missed errors.
@@ 4.1 Experimental Setup
We compare SummQ against strong baselines from three categories: (1) Supervised Fine-Tuning: TextRank (Mihalcea & Tarau, 2004; Jeong et al., 2025), LongT5 (Guo et al., 2021), U.former (Bertsch et al., 2023), SLED (Ivgi et al., 2023), and CachED (Saxena et al., 2025). (2) Prompting: Proprietary LLMs including GPT-4o, GPT-4.1, GPT-5, o3, and open-source models such as DeepSeek-R1 and Qwen3-32B. (3) Multi-Agent Systems: HM-SR (Jeong et al., 2025) and C.MultiLLM (Fang et al., 2024).
@@
In our experiments, we consider two configurations of our approach: SummQsolo, where each component employs a single agent for efficient and straightforward deployment, and SummQcombo, where each component leverages multiple agents in an ensemble manner to facilitate collaborative generation and review. By default, we use GPT-4o as the agent backbone for both SummQsolo and SummQcombo, deploying three agents in each component for SummQcombo unless otherwise specified. Moreover, the number of iterations $T_{\text{iter}}$ is set to three for all configurations and the number of debate rounds $T_{\text{debate}}$ is set to one. Prompts used in SummQ are in Appendix F.
@@ 4.3 Human Evaluation Results
We conduct a human evaluation using 20 randomly selected NLP papers published after June 2024, with five Ph.D. students as judges. Each judge compares two summaries considering Informativeness, Coherence, and Factuality. The performance is measured by the winning rate. To address length bias, we include both SummQcombo and a rephrased version SummQcomboR with shortened summaries. Details on the evaluation protocol and selected papers are in Appendix D.
@@ 5 Analysis
We divide each document into five equal segments and use GPT-4.1 to map quiz questions to segments. As shown in Figure 7, quiz questions initially focus on the beginning and end, which aligns well with the findings of Liu et al. (2024), but coverage becomes more balanced by iteration 3, with increased attention to middle segments. This shift indicates a more holistic document understanding, which is crucial for generating summaries that accurately reflect the entire content.
@@ Acknowledgement
This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
@@ Appendix C Accuracy of Questions Evolve
Appendix D Human Evaluation
@@
Appendix E Case Study
@@
Appendix F Prompts
F.1 Prompts of Summary Generators
@@
F.2 Prompts of Quiz Generators
@@
F.3 Prompts of Summary Reviewers
@@
F.4 Prompts of Quiz Reviewers

### v3

@@ Abstract
Long document summarization remains a significant challenge for current large language models (LLMs), as existing approaches commonly struggle with information loss, factual inconsistencies, and coherence issues when processing excessively long documents. We propose SummQ, a novel adversarial multi-agent framework that addresses these limitations through collaborative intelligence between specialized agents operating in two complementary domains: summarization and quizzing. Our approach employs summary generators and reviewers that work collaboratively to create and evaluate comprehensive summaries, while quiz generators and reviewers create comprehension questions that serve as continuous quality checks for the summarization process. This adversarial dynamic, enhanced by an examinee agent that validates whether the generated summary contains the information needed to answer the quiz questions, enables iterative refinement through multifaceted feedback mechanisms. We evaluate SummQ on three widely used long document summarization benchmarks. Experimental results demonstrate that our framework significantly outperforms existing state-of-the-art methods across ROUGE and BERTScore metrics, as well as in LLM-as-a-Judge and human evaluations. Our comprehensive analyses reveal the effectiveness of the multi-agent collaboration dynamics, the influence of different agent configurations, and the impact of the quizzing mechanism. This work establishes a new approach for long document summarization that uses adversarial agentic collaboration to improve summarization quality. In addition, we release the code to foster research along this line. [^1] https://github.com/weixuan-wang123/SummQ
@@ 1 Introduction
Summarization has become increasingly critical in modern natural language processing, as organizations and individuals face an ever-growing volume of textual information that requires efficient processing and comprehension (DBLP:journals/air/GambhirG17; zhao2020seal; zhang2021summ). Prior works in summarization have primarily focused on short to medium-length documents, where models can effectively capture the essential content and generate coherent summaries (see-etal-2017-get; fabbri-etal-2019-multi). Recently, there has been an increasing interest in long document document summarization, driven by the need to process extensive texts such as research articles, legal documents, and books (huang-etal-2021-efficient; kryscinski-etal-2022-booksum; saxena-keller-2024-select).
Recent large language models (LLMs) have shown promising results for summarization tasks (pu2023summarization; laban-etal-2023-summedits; keswani2024abstractive). However, existing methods struggle with long documents, often leading to significant information loss, factual inconsistencies, and difficulty maintaining coherence across lengthy texts (DBLP:journals/csur/KohJLP23). Current approaches often fail to capture the nuanced relationships between distant parts of a document, resulting in summaries that may miss crucial information or introduce hallucinations (chrysostomou-etal-2024-investigating; tang-etal-2024-tofueval; xia-etal-2024-hallucination). Recently, multi-agent systems has demonstrated potential for improving complex reasoning tasks through collaborative interactions (DBLP:conf/ijcai/GuoCWCPCW024), yet their application to long document summarization remains underexplored (fang2024multi; kim-kim-2025-nexussum).
@@
To validate the effectiveness of SummQ, we conduct extensive experiments on three long document summarization tasks including MENSA (saxena-keller-2024-select), BookSum (kryscinski-etal-2022-booksum), and GovReport (huang-etal-2021-efficient). Our results demonstrate that SummQ significantly outperforms existing state-of-the-art methods in terms of ROUGE (lin-2004-rouge) and BERTScore (DBLP:conf/iclr/ZhangKWWA20), as well as LLM-as-a-Judge and human evaluations. Furthermore, more in-depth analyses highlight the effectiveness of our multi-agent framework in enhancing summary quality, the coverage of the generated quizzes, and the impact of various agent configurations.
@@ 2 Related Work
Recent advances in LLMs have enabled the development of multi-agent systems that harness the strengths of multiple agents to tackle complex tasks (wang2024survey; DBLP:conf/ijcai/GuoCWCPCW024; xi2025rise). These systems typically involve agent collaboration to boost performance, as seen in multi-agent debating (DBLP:conf/icml/Du00TM24; DBLP:conf/emnlp/XiongD00023; DBLP:journals/corr/abs-2310-20151; DBLP:conf/acl/TangZ0L0ZCG24) and discussion (DBLP:conf/acl/ChenSB24; DBLP:conf/naacl/SahaLCBWL24) for reasoning over short texts (DBLP:conf/icml/Du00TM24; DBLP:conf/acl/TangZ0L0ZCG24), paper review (DBLP:journals/corr/abs-2311-08152), dataset synthesis (wang2025odysseybench), machine translation (DBLP:journals/corr/abs-2405-11804), and code generation (DBLP:journals/corr/abs-2312-13010; DBLP:conf/coling/WangR0LBCYZYSL25). Collaboration among agents introduces diverse perspectives and complementary skills, leading to higher-quality and more robust outputs.
@@
Long document summarization has seen various methods, including architectural optimizations (e.g., sparse attention (DBLP:conf/emnlp/0006ZWX0Y21; ivgi2023efficient; bertsch2023unlimiformer), long-context finetuning (DBLP:journals/corr/abs-2004-05150; guo2021longt5), memory augmentation (cui2021sliding; saxena2025end), window extension (press2021train; DBLP:journals/corr/abs-2306-15595; DBLP:journals/ijon/SuALPBL24; DBLP:conf/acl/YenG024)) and chunking strategies (e.g., sliding window (zaheer2020big; DBLP:conf/eacl/PangNKSZX23). LLMs with improved long-context abilities have shifted the field toward leveraging their strong language skills for summarization (goyal2022news; DBLP:conf/acl/RatnerLBRMAKSLS23; keswani2024abstractive), but still face challenges with context limits and maintaining coherence (pu2023summarization; liu2023lost). Multi-agent systems have been explored to address these issues, enabling collaborative, more accurate summaries (DBLP:journals/corr/abs-2402-11550; fang2024multi; jeong2025agent), though many still rely on self-verification, leading to biases and missed errors.
@@ 4.1 Experimental Setup
We compare SummQ against strong baselines from three categories: (1) Supervised Fine-Tuning: TextRank (mihalcea2004textrank; jeong2025agent), LongT5 (guo2021longt5), U.former (bertsch2023unlimiformer), SLED (ivgi2023efficient), and CachED (saxena2025end). (2) Prompting: Proprietary LLMs including GPT-4o, GPT-4.1, GPT-5, o3, and open-source models such as DeepSeek-R1 and Qwen3-32B. (3) Multi-Agent Systems: HM-SR (jeong2025agent) and C.MultiLLM (fang2024multi).
@@
In our experiments, we consider two configurations of our approach: SummQsolo, where each component employs a single agent for efficient and straightforward deployment, and SummQcombo, where each component leverages multiple agents in an ensemble manner to facilitate collaborative generation and review. By default, we use GPT-4o as the agent backbone for both SummQsolo and SummQcombo, deploying three agents in each component for SummQcombo unless otherwise specified. Moreover, the number of iterations $T_{\text{iter}}$ is set to three for all configurations and the number of debate rounds $T_{\text{debate}}$ is set to one. Prompts used in SummQ are in Appendix G.
@@ 4.3 Human Evaluation Results
We conduct a human evaluation using 20 randomly selected NLP papers published after June 2024, with five Ph.D. students as judges. Each judge compares two summaries considering Informativeness, Coherence, and Factuality. The performance is measured by the winning rate. To address length bias, we include both SummQcombo and a rephrased version SummQcomboR with shortened summaries. Details on the evaluation protocol and selected papers are in Appendix E.
@@ 5 Analysis
We divide each document into five equal segments and use GPT-4.1 to map quiz questions to segments. As shown in Figure 7, quiz questions initially focus on the beginning and end, which aligns well with the findings of liu-etal-2024-lost, but coverage becomes more balanced by iteration 3, with increased attention to middle segments. This shift indicates a more holistic document understanding, which is crucial for generating summaries that accurately reflect the entire content.
@@ 6 Conclusion
Ethics Statement
This work introduces SummQ, an adversarial multi-agent framework for long document summarization using large language models (LLMs). All experiments were conducted using publicly available datasets and LLMs, strictly adhering to their respective licenses and usage policies. No human subjects were involved in the development or evaluation of the system, except for the human evaluation, which was performed by consenting Ph.D. students with relevant expertise. We acknowledge that LLMs and datasets may contain inherent biases, which could affect the generated summaries and quiz questions. We encourage responsible use of our framework, with attention to fairness, transparency, and accountability in downstream applications.
Reproducibility Statement
We are committed to reproducibility in this work. Detailed descriptions of the SummQ framework, including algorithms, agent configurations, and collaboration mechanisms, are provided in Section 3. Experimental setups, including model backbones, datasets, evaluation metrics, and baseline comparisons, are thoroughly described in Section 4. All datasets used are standard benchmarks, and references are included for accessibility. Prompts and implementation details are provided in Appendix. To further support reproducibility, we will release our code and experiment scripts upon publication, enabling other researchers to replicate and extend our results.
The Use of Large Language Models (LLMs)
In preparing this work, we utilize large language models (LLMs) as general-purpose tools to assist with writing polish and grammar correction. The LLMs are not involved in research ideation, experimental design, or substantive content generation. Their role is limited to improving the clarity and readability of the text, ensuring grammatical accuracy, and refining the presentation of our findings. All scientific contributions, analyses, and conclusions are solely the work of the authors.
@@ Acknowledgement
This project has received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant numbers 10039436].
@@ Appendix C Accuracy of Questions Evolve
Appendix D Cost Analysis
The cost analysis in Table 5 reveals significant differences in computational resource requirements across different approaches. While simple prompting baseline maintains the lowest cost at \$0.18 per instance with minimal input and output token usage, SummQsolo demonstrates a moderate increase in resource consumption, requiring \$1.95 per instance with 0.43M input tokens and 6.97K output tokens. This reflects the additional computational overhead of our iterative refinement process compared to baseline approaches. The SummQcombo variant shows the highest resource requirements at \$14.45 per instance, consuming 3.30M input tokens and generating 24.96K output tokens, which is attributable to the collaborative multi-agent framework involving multiple summary and quiz generators, reviewers, and the iterative debate process. Despite the higher computational cost, the substantial improvements in summary quality and quiz generation accuracy demonstrated throughout our evaluation justify this investment, particularly for applications where high-quality outputs are prioritized over computational efficiency.
Table 5: Average token usage and cost (in USD) per example of different methods on MENSA with GPT-4o.
Input Tokens Output Tokens Cost (USD)
Prompting 0.04M 00.23K 00.18
SummQsolo 0.43M 06.97K 01.95
SummQcombo 3.30M 24.96K 14.45
Appendix E Human Evaluation
@@
Appendix F Case Study
@@
Appendix G Prompts
G.1 Prompts of Summary Generators
@@
G.2 Prompts of Quiz Generators
@@
G.3 Prompts of Summary Reviewers
@@
G.4 Prompts of Quiz Reviewers